# An Overview of Systems-Theoretic Guarantees in Data-Driven Model Predictive Control

### v1

@@ 1 INTRODUCTION
As an alternative to indirect approaches, designing controllers directly from data without intermediate system identification has received increasing attention in the recent literature [3]. The Fundamental Lemma by Willems et al. [4] provides a unifying framework for direct data-driven control. It states that one persistently exciting trajectory of an unknown linear time-invariant (LTI) system can be used to parametrize all other trajectories. This allows to address various analysis and controller design problems, that are traditionally solved using model knowledge, based purely on data, see [5] for a recent survey. One of the most promising applications of the Fundamental Lemma is the design of data-driven model predictive control (MPC) schemes. MPC is a powerful modern control technique which can handle constraints, performance criteria, and nonlinear dynamics [6]. It relies on a receding-horizon principle: At each time step, an open-loop optimal control problem is solved, the first component of the optimal input sequence is applied to the plant, and the process is repeated at the next time step using a new measurement. While classical MPC schemes assume precise model knowledge, they can be enhanced via data using learning-based [7] and adaptive [8, 9, 10] approaches.
@@
Notation: We denote the set of nonnegative integers by $\mathbb{I}_{\geq 0}$ and integers in the interval $[a,b]$ by $\mathbb{I}_{[a,b]}$. For a sequence $\{\mathbf{u}_{k}\}_{k=0}^{N-1}$, we define the stacked column vector $\mathbf{u}_{[k_{1},k_{2}]}\coloneqq\begin{bmatrix}\mathbf{u}_{k_{1}}^{\top}&\mathbf{u}_{k_{1}+1}^{\top}&\dots&\mathbf{u}_{k_{2}}^{\top}\end{bmatrix}^{\top}$ and we abbreviate $\mathbf{u}\coloneqq\mathbf{u}_{[0,N-1]}$ for the vector containing the full sequence. Further, we define the Hankel matrix
@@
For a vector $\mathbf{x}$, we denote the $p$-norm by $\lVert\mathbf{x}\rVert_{p}$ and we define the weighted norm $\lVert\mathbf{x}\rVert_{\mathbf{P}}=\sqrt{\mathbf{x}^{\top}\mathbf{P}\mathbf{x}}$ for some matrix $\mathbf{P}=\mathbf{P}^{\top}$.
@@ 2.1 The Fundamental Lemma
with state $\mathbf{x}_{k}\in\mathbb{R}^{n}$, input $\mathbf{u}_{k}\in\mathbb{R}^{m}$, output $\mathbf{y}_{k}\in\mathbb{R}^{p}$, and time $k\in\mathbb{I}_{\geq 0}$. We make the standing assumption that $(\mathbf{A},\mathbf{B})$ is controllable and $(\mathbf{A},\mathbf{C})$ is observable. The matrices $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, $\mathbf{D}$, are unknown but, instead, an input-output data trajectory $\{\mathbf{u}_{k}^{\mathrm{d}},\mathbf{y}_{k}^{\mathrm{d}}\}_{k=0}^{N-1}$ is available. We require the corresponding input signal to be sufficiently rich in the following sense [4].
@@ 3.1 Data-driven MPC with noise-free data
The Fundamental Lemma can be used to set up data-driven MPC schemes by replacing the commonly used state-space model (compare Section 2.2) via data-dependent Hankel matrices. Let us consider the control goal of tracking an input-output setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ while satisfying constraints $\mathbf{u}_{k}\in\mathbb{U}$ and $\mathbf{y}_{k}\in\mathbb{Y}$ for all $k\in\mathbb{I}_{\geq 0}$ with sets $\mathbb{U}\subseteq\mathbb{R}^{m}$, $\mathbb{Y}\subseteq\mathbb{R}^{p}$. The Fundamental Lemma allows to achieve these two objectives based only on data $\{\mathbf{u}_{k}^{\mathrm{d}},\mathbf{y}_{k}^{\mathrm{d}}\}_{k=0}^{N-1}$ by solving, at time $t$, the following optimal control problem
@@
The input-output trajectory predicted at time $t$ is denoted by $\mathbf{\bar{u}}(t)=\mathbf{\bar{u}}_{[-n,L-1]}(t)\in\mathbb{U}^{L+n}$, $\mathbf{\bar{y}}(t)=\mathbf{\bar{y}}_{[-n,L-1]}(t)\in\mathbb{Y}^{L+n}$. The constraint in Equation 4b is based on the Fundamental Lemma and ensures that $\{\mathbf{\bar{u}}_{k}(t),\mathbf{\bar{y}}_{k}(t)\}_{k=-n}^{L-1}$ is a trajectory of the LTI system. The cost penalizes the difference of the predicted trajectory w.r.t. the setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ for user-specified positive definite matrices ${\mathbf{Q}}$, ${\mathbf{R}}$. Note that the predicted trajectory is of length $L+n$. In the constraint 4c, the first $n$ components $\{\mathbf{\bar{u}}_{k}(t),\mathbf{\bar{y}}_{k}(t)\}_{k=-n}^{-1}$ are used to initialize the prediction based on the most recent $n$ input-output measurements $\{\mathbf{u}_{k},\mathbf{y}_{k}\}_{k=t-n}^{t-1}$. This initialization over $n$ time steps is required when using an input-output prediction model to ensure that the internal states of the prediction and the plant coincide, i.e., to implicitly enforce a constraint analogous to the initial condition in Equation 3b. Instead of the system order $n$, any upper bound can be used (to be precise, an upper bound on the lag is sufficient). Further, Equation 4d contains the input-output constraints. Note that, for convex polytopic constraint sets $\mathbb{U}$, $\mathbb{Y}$, the optimization problem 4 is a convex quadratic program. We denote the optimal solution of the optimization problem 4 by $\alpha^{*}(t)$, $\mathbf{\bar{u}}^{*}(t)$, $\mathbf{\bar{y}}^{*}(t)$. It can be used to set up a data-driven MPC scheme in a receding-horizon fashion analogous to the model-based case in Section 2.2, see Algorithm 3.1.
@@
Algorithm 3.1 allows to control unknown LTI systems based only on input-output data and without explicit model knowledge. The main difference to model-based MPC as in Section 2.2 is the usage of data-dependent Hankel matrices for prediction, instead of a state-space model. If the data are PE, the optimization problem 4 returns the same optimal input as using a state-space model. Since the Fundamental Lemma directly parametrizes input-output trajectories, Algorithm 3.1 inherently is an output-feedback MPC scheme. In contrast, model-based output-feedback MPC schemes typically involve an observer whose estimation error needs to be accounted for.
@@
This implicitly ensures that the internal state corresponding to the input-output trajectory $\mathbf{\bar{u}}(t)$, $\mathbf{\bar{y}}(t)$ is equal to the steady-state $\mathbf{x}^{\mathrm{s}}$ corresponding to $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ at time $L$. With this modification, it can be shown under mild assumptions (e.g., PE) that, if the optimization problem is feasible at initial time $t=0$, then it is recursively feasible for all $t\in\mathbb{I}_{\geq 0}$ and Algorithm 3.1 exponentially stabilizes $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ in closed loop [51]. While terminal equality constraints are simple to implement and to study theoretically, they have significant drawbacks in terms of robustness and the size of the region of attraction. In particular, the latter only includes points from which the optimization problem is initially feasible, i.e., from which the setpoint can be reached within $L$ time steps while satisfying the constraints.
@@
All approaches discussed so far require that the setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ is a feasible equilibrium for the unknown system, i.e., there exists $\mathbf{x}^{\mathrm{s}}\in\mathbb{R}^{n}$ such that $\mathbf{x}^{\mathrm{s}}=\mathbf{A}\mathbf{x}^{\mathrm{s}}+\mathbf{B}\mathbf{u}^{\mathrm{s}},\>\>\mathbf{y}^{\mathrm{s}}=\mathbf{C}\mathbf{x}^{\mathrm{s}}+\mathbf{D}\mathbf{u}^{\mathrm{s}}$. This condition is non-trivial to verify, especially when no model is available, and therefore it can be a critical limitation for practical applications of data-driven MPC. A powerful way to resolve this problem is to optimize over the equilibrium $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ online and penalize its distance to the actual target setpoint in the cost. Such an artificial equilibrium has been proposed in model-based MPC [61] and it leads to strong theoretical results as well as useful practical methods. In particular, data-driven MPC with artificial equilibrium guarantees exponential stability of the unknown optimal reachable equilibrium for a given setpoint and cost function [62]. This includes, e.g., the case that only an output setpoint without corresponding input component is available. Additional advantages include theoretical guarantees for online setpoint changes, improved robustness, and a significantly larger region of attraction compared to a standard MPC scheme with terminal equality constraints. Further, online optimization over an artificial equilibrium plays a crucial role when using data-driven MPC to control unknown nonlinear systems, see Section 5 for details.
To summarize, if the data are noise-free, then the Fundamental Lemma allows to predict future trajectories exactly such that, under appropriate assumptions (long prediction horizon) or modifications (terminal ingredients), closed-loop stability under data-driven MPC can be guaranteed. The main technical challenge in the analysis is that the cost function penalizes input-output values due to the employed input-output prediction model. This only implies a positive semidefinite cost in the state, in contrast to model-based MPC which commonly assumes a positive definite cost, compare the optimization problem 3. Therefore, the analysis of data-driven MPC requires additional detectability arguments [63].
@@ 4.1 Data-driven MPC with noisy data
Suppose that both the offline data used for prediction via the Fundamental Lemma as well as the online data used to specify initial conditions at time $t$ are affected by output measurement noise. To be precise, instead of the exact output values, we have access to
@@
compare Figure 1. Further, the noise is bounded by some $\bar{\mathbf{\varepsilon}}>0$, i.e., $\lVert\mathbf{\varepsilon}_{k}^{\mathrm{d}}\rVert_{\infty}\leq\bar{\mathbf{\varepsilon}}$ for $k\in\mathbb{I}_{[0,N-1]}$ and $\lVert\mathbf{\varepsilon}_{t}\rVert_{\infty}\leq\bar{\mathbf{\varepsilon}}$ for $t\in\mathbb{I}_{\geq 0}$. The following optimization problem provides the basis for controlling unknown linear systems based on noisy input-output data.
@@ 4.1.1 Closed-loop guarantees for bounded noise
Similar to Section 3.1, Algorithm 4.1 does not guarantee closed-loop stability unless the prediction horizon is sufficiently long or terminal ingredients are added. It was shown in [51] that stability and robustness can be ensured by adding terminal equality constraints as in Equation 5 to the optimization problem 8. In this case, due to a technical controllability argument, proving theoretical guarantees requires the application in an $n$-step fashion [67]. This means that, in each MPC iteration, the first $n$ optimal input components $\mathbf{\bar{u}}_{[0,n-1]}^{*}(t)$ are applied to the system before repeating the optimization. With these modifications, robust data-driven MPC with terminal equality constraints guarantees that the internal state $\mathbf{x}_{t}$ converges exponentially to a neighborhood of the steady-state $\mathbf{x}^{\mathrm{s}}$ corresponding to $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$, where the size of the neighborhood depends on the noise level. Mathematically, there exist $10<c<$ 1 as well as a continuous and strictly increasing function $\beta$ with $\beta(0)=0$ such that
@@
compare [51, Theorem 3]. For $\bar{\mathbf{\varepsilon}}=0$, the closed loop converges to the desired setpoint. These theoretical guarantees require analogous assumptions as in the noise-free case (e.g., PE data, $n$ is an upper bound on the system order, compare Section 3.1) and additionally that the noise level $\bar{\mathbf{\varepsilon}}$ is sufficiently small and that the regularization parameters scale as $\lambda_{\mathbf{\alpha}}\sim\bar{\mathbf{\varepsilon}}$, $\lambda_{\mathbf{\sigma}}\sim\frac{1}{\bar{\mathbf{\varepsilon}}}$. The latter conditions ensure that, in the noise-free case $\bar{\mathbf{\varepsilon}}=0$, robust data-driven MPC (Algorithm 4.1) reduces to the nominal scheme (Algorithm 3.1) since $\lambda_{\mathbf{\alpha}}=0$ and $\lambda_{\mathbf{\sigma}}\to\infty$ enforces $\mathbf{\sigma}(t)=0$. Beyond this stability result, the theoretical analysis reveals direct connections between the data quality and the closed-loop performance. In particular, the asymptotic tracking error decreases and the region of attraction increases if the minimum singular value of the input Hankel matrix $\mathbf{H}_{L+2n}(u)$ increases, which corresponds to a quantitative PE condition [39]. The possiblity to establish such end-to-end insights is a valuable feature of direct data-driven control methods, which is often harder to achieve in indirect approaches. While the original paper [51] required an additional non-convex constraint $\lVert\mathbf{\sigma}(t)\rVert_{\infty}\leq\bar{\mathbf{\varepsilon}}(1+\lVert\mathbf{\alpha}(t)\rVert_{1})$ to bound the slack variable, it was later shown in [58] that this constraint can be dropped when $\lambda_{\mathbf{\sigma}}\sim\frac{1}{\bar{\mathbf{\varepsilon}}}$, see also [68, Theorem 4.1].
@@
As in the noise-free case, the occurrence of terminal equality constraints poses practical limitations, e.g., a small region of attraction, and alternative means of guaranteeing closed-loop stability are available. Roughly speaking, any model-based MPC scheme which admits inherent robustness properties against disturbances can be transformed into a data-driven MPC scheme with closed-loop guarantees in the presence of noise [69]. Beyond this general principle, the literature contains various concrete results on closed-loop guarantees in robust data-driven MPC. For example, in the noisy data case, terminal ingredients can be designed based on robust data-driven output-feedback design [55]. By combining continuity of data-driven MPC w.r.t. noise [70] and inherent robustness properties [69, 71], one can show that the resulting data-driven MPC scheme admits closed-loop guarantees also in the presence of noise, see [68, Section 4.5.2] for a discussion. Likewise, data-driven MPC without terminal ingredients practically exponentially stabilizes the closed-loop system, assuming that the prediction horizon $L$ is sufficiently long [58]. Notably, in both cases (terminal ingredients and $L$ sufficiently long), stability and robustness can be proven even for a $1$-step MPC scheme as in Algorithm 4.1. Alternatively, one can optimize over an artifical equilibrium online which provides powerful practical features (no knowledge required if $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ is a feasible equilibrium, improved robustness, increased region of attraction, handling online setpoint changes) and also admits closed-loop guarantees in the presence of noise [70].

### v2

@@ 1 INTRODUCTION
As an alternative to indirect approaches, designing controllers directly from data without intermediate system identification has received increasing attention in the recent literature [3]. The Fundamental Lemma by Willems et al. [4] provides a unifying framework for direct data-driven control. It states that one persistently exciting trajectory of an unknown linear time-invariant (LTI) system can be used to parametrize all other trajectories. This allows one to address various analysis and controller design problems, that are traditionally solved using model knowledge, based purely on data, see [5] for a recent survey. One of the most promising applications of the Fundamental Lemma is the design of data-driven model predictive control (MPC) schemes. MPC is a powerful modern control technique which can handle constraints, performance criteria, and nonlinear dynamics [6]. It relies on a receding-horizon principle: At each time step, an open-loop optimal control problem is solved, the first component of the optimal input sequence is applied to the plant, and the process is repeated at the next time step using a new measurement. While classical MPC schemes assume precise model knowledge, they can be enhanced via data using learning-based [7] and adaptive [8, 9, 10] approaches.
@@
Notation: We denote the set of nonnegative integers by $\mathbb{I}_{\geq 0}$ and integers in the interval $[a,b]$ by $\mathbb{I}_{[a,b]}$. For a vector $\mathbf{x}$, we denote the $p$-norm by $\lVert\mathbf{x}\rVert_{p}$ and we define the weighted norm $\lVert\mathbf{x}\rVert_{\mathbf{P}}=\sqrt{\mathbf{x}^{\top}\mathbf{P}\mathbf{x}}$ for some matrix $\mathbf{P}=\mathbf{P}^{\top}$. For a sequence $\{\mathbf{u}_{k}\}_{k=0}^{N-1}$, we define the stacked column vector $\mathbf{u}_{[k_{1},k_{2}]}\coloneqq\begin{bmatrix}\mathbf{u}_{k_{1}}^{\top}&\mathbf{u}_{k_{1}+1}^{\top}&\dots&\mathbf{u}_{k_{2}}^{\top}\end{bmatrix}^{\top}$ and we abbreviate $\mathbf{u}\coloneqq\mathbf{u}_{[0,N-1]}$ for the vector containing the full sequence. Further, we define the Hankel matrix
@@ 2.1 The Fundamental Lemma
with state $\mathbf{x}_{k}\in\mathbb{R}^{n}$, input $\mathbf{u}_{k}\in\mathbb{R}^{m}$, output $\mathbf{y}_{k}\in\mathbb{R}^{p}$, and time $k\in\mathbb{I}_{\geq 0}$. We make the standing assumption that $(\mathbf{A},\mathbf{B})$ is controllable and $(\mathbf{A},\mathbf{C})$ is observable. The matrices $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, $\mathbf{D}$, are unknown but an input-output data trajectory $\{\mathbf{u}_{k}^{\mathrm{d}},\mathbf{y}_{k}^{\mathrm{d}}\}_{k=0}^{N-1}$ is available. We require the corresponding input signal to be sufficiently rich in the following sense [4].
@@ 3.1 Data-driven MPC with noise-free data
The Fundamental Lemma can be used to set up data-driven MPC schemes by replacing the commonly used state-space model (compare Section 2.2) via data-dependent Hankel matrices. Let us consider the control goal of tracking an input-output setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ while satisfying constraints $\mathbf{u}_{k}\in\mathbb{U}$ and $\mathbf{y}_{k}\in\mathbb{Y}$ for all $k\in\mathbb{I}_{\geq 0}$ with sets $\mathbb{U}\subseteq\mathbb{R}^{m}$, $\mathbb{Y}\subseteq\mathbb{R}^{p}$. In the following, we assume that the setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ is feasible for the system dynamics 1 and we discuss possibilities for relaxing this assumption below. The Fundamental Lemma can be used to design an MPC scheme based only on data $\{\mathbf{u}_{k}^{\mathrm{d}},\mathbf{y}_{k}^{\mathrm{d}}\}_{k=0}^{N-1}$ by solving, at time $t$, the following optimal control problem
@@
The input-output trajectory predicted at time $t$ is denoted by $\mathbf{\bar{u}}(t)=\mathbf{\bar{u}}_{[-n,L-1]}(t)\in\mathbb{U}^{L+n}$, $\mathbf{\bar{y}}(t)=\mathbf{\bar{y}}_{[-n,L-1]}(t)\in\mathbb{Y}^{L+n}$. The constraint in Equation 4b is based on the Fundamental Lemma and ensures that $\{\mathbf{\bar{u}}_{k}(t),\mathbf{\bar{y}}_{k}(t)\}_{k=-n}^{L-1}$ is a trajectory of the LTI system. The cost penalizes the difference of the predicted trajectory with respect to the setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ for user-specified positive definite matrices $\mathbf{Q}$}, $\mathbf{R}$}. Note that the predicted trajectory is of length $L+n$. In the constraint 4c, the first $n$ components $\{\mathbf{\bar{u}}_{k}(t),\mathbf{\bar{y}}_{k}(t)\}_{k=-n}^{-1}$ are used to initialize the prediction based on the most recent $n$ input-output measurements $\{\mathbf{u}_{k},\mathbf{y}_{k}\}_{k=t-n}^{t-1}$. This initialization over $n$ time steps is required when using an input-output prediction model to ensure that the internal states of the prediction and the plant coincide, i.e., to implicitly enforce a constraint analogous to the initial condition in Equation 3b. Instead of the system order $n$, any upper bound can be used (to be precise, an upper bound on the lag is sufficient). Further, Equation 4d contains the input-output constraints. Note that, for convex polytopic constraint sets $\mathbb{U}$, $\mathbb{Y}$, the optimization problem 4 is a convex quadratic program. We denote the optimal solution of the optimization problem 4 by $\alpha^{*}(t)$, $\mathbf{\bar{u}}^{*}(t)$, $\mathbf{\bar{y}}^{*}(t)$. It can be used to set up a data-driven MPC scheme in a receding-horizon fashion analogous to the model-based case in Section 2.2, see Algorithm 3.1.
@@
Algorithm 3.1 can be used to control unknown LTI systems based only on input-output data and without explicit model knowledge. The main difference to model-based MPC in Section 2.2 is the usage of data-dependent Hankel matrices for prediction, instead of a state-space model. If the data are PE, the optimization problem 4 returns the same optimal input as using a state-space model. Since the Fundamental Lemma directly parametrizes input-output trajectories, Algorithm 3.1 inherently is an output-feedback MPC scheme. In contrast, model-based output-feedback MPC schemes typically involve an observer.
@@
This ensures that the internal state corresponding to the input-output trajectory $\mathbf{\bar{u}}(t)$, $\mathbf{\bar{y}}(t)$ is equal to the steady-state $\mathbf{x}^{\mathrm{s}}$ corresponding to $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ at time $L$. With this modification, it can be shown under mild assumptions (e.g., PE) that, if the optimization problem is feasible at initial time $t=0$, then it is recursively feasible for all $t\in\mathbb{I}_{\geq 0}$ and Algorithm 3.1 exponentially stabilizes $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ in closed loop [51]. While terminal equality constraints are simple to implement and to study theoretically, they have significant drawbacks in terms of robustness and the size of the region of attraction. In particular, the latter only includes points with respect to which the optimization problem is initially feasible, i.e., from which the setpoint can be reached within $L$ time steps while satisfying the constraints.
@@
All approaches discussed so far require that the setpoint $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ is a feasible equilibrium for the unknown system, i.e., there exists $\mathbf{x}^{\mathrm{s}}\in\mathbb{R}^{n}$ such that $\mathbf{x}^{\mathrm{s}}=\mathbf{A}\mathbf{x}^{\mathrm{s}}+\mathbf{B}\mathbf{u}^{\mathrm{s}},\>\>\mathbf{y}^{\mathrm{s}}=\mathbf{C}\mathbf{x}^{\mathrm{s}}+\mathbf{D}\mathbf{u}^{\mathrm{s}}$. If this is not the case, then, for example, the terminal equality constraint 5 cannot be feasible. Feasibility of $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ is non-trivial to verify, especially when no model is available, and therefore it can be a critical limitation for practical applications of data-driven MPC. A powerful way to resolve this problem is to optimize over the equilibrium $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ online and penalize its distance to the actual target setpoint in the cost. Such an artificial equilibrium has been proposed in model-based MPC [61] and it leads to strong theoretical results as well as useful practical methods. In particular, data-driven MPC with artificial equilibrium guarantees exponential stability of the unknown optimal reachable equilibrium for a given setpoint and cost function [62]. This includes, e.g., the case that only an output setpoint without corresponding input component is available. Additional advantages include theoretical guarantees for online setpoint changes, improved robustness, and a significantly larger region of attraction compared to a standard MPC scheme with terminal equality constraints. Further, time-varying reference trajectories can be handled as well by translating corresponding model-based MPC results [63]. Finally, online optimization over an artificial equilibrium plays a crucial role when using data-driven MPC to control unknown nonlinear systems, see Section 5 for details.
To summarize, if the data are noise-free, then the Fundamental Lemma can be used to predict future trajectories exactly such that, under appropriate assumptions (long prediction horizon) or modifications (terminal ingredients), closed-loop stability under data-driven MPC can be guaranteed. The main technical challenge in the analysis is that the cost function penalizes input-output values due to the employed input-output prediction model. This only implies a positive semidefinite cost in the state, in contrast to model-based MPC which commonly assumes a positive definite cost, compare the optimization problem 3. Therefore, the analysis of data-driven MPC requires additional detectability arguments [64].
@@ 4.1 Data-driven MPC with noisy data
Suppose that both the offline data used for prediction via the Fundamental Lemma as well as the online data used to specify initial conditions at time $t$ are affected by output measurement noise (disturbances can be handled similarly [67]). To be precise, instead of the exact output values, we have access to
@@
compare Figure 1. Further, the noise is bounded by some $\bar{\mathbf{\varepsilon}}>0$ in the sense that $\lVert\mathbf{\varepsilon}_{k}^{\mathrm{d}}\rVert_{\infty}\leq\bar{\mathbf{\varepsilon}}$ for $k\in\mathbb{I}_{[0,N-1]}$ and $\lVert\mathbf{\varepsilon}_{t}\rVert_{\infty}\leq\bar{\mathbf{\varepsilon}}$ for $t\in\mathbb{I}_{\geq 0}$. The following optimization problem provides the basis for controlling unknown linear systems based on noisy input-output data.
@@ 4.1.1 Closed-loop guarantees for bounded noise
Similar to Section 3.1, Algorithm 4.1 does not guarantee closed-loop stability unless the prediction horizon is sufficiently long or terminal ingredients are added. It was shown in [51] that stability and robustness can be ensured by adding terminal equality constraints as in Equation 5 to the optimization problem 8. In this case, due to a technical controllability argument, proving theoretical guarantees requires the application in an $n$-step fashion [69]. This means that, in each MPC iteration, the first $n$ optimal input components $\mathbf{\bar{u}}_{[0,n-1]}^{*}(t)$ are applied to the system before repeating the optimization. With these modifications, robust data-driven MPC with terminal equality constraints guarantees that the internal state $\mathbf{x}_{t}$ converges exponentially to a neighborhood of the steady-state $\mathbf{x}^{\mathrm{s}}$ corresponding to $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$, where the size of the neighborhood depends on the noise level. Mathematically, there exist $0<c<1$ as well as a continuous and strictly increasing function $\beta$ with $\beta(0)=0$ such that
@@
compare [51, Theorem 3]. For $\bar{\mathbf{\varepsilon}}=0$, the closed loop converges to the desired setpoint as in the noise-free case of Section 3.1. These theoretical guarantees require analogous assumptions as in the noise-free case (e.g., PE data, $n$ is an upper bound on the system order, compare Section 3.1) and additionally that the noise level $\bar{\mathbf{\varepsilon}}$ is sufficiently small and that the regularization parameters scale as $\lambda_{\mathbf{\alpha}}\sim\bar{\mathbf{\varepsilon}}$, $\lambda_{\mathbf{\sigma}}\sim\frac{1}{\bar{\mathbf{\varepsilon}}}$. The latter conditions ensure that, in the noise-free case $\bar{\mathbf{\varepsilon}}=0$, robust data-driven MPC (Algorithm 4.1) reduces to the nominal scheme (Algorithm 3.1) since $\lambda_{\mathbf{\alpha}}=0$ and $\lambda_{\mathbf{\sigma}}\to\infty$ enforces $\mathbf{\sigma}(t)=0$. Beyond this stability result, the theoretical analysis reveals direct connections between the data quality and the closed-loop performance: The asymptotic tracking error decreases and the region of attraction increases if the minimum singular value of the input Hankel matrix $\mathbf{H}_{L+2n}(u)$ increases, which corresponds to a quantitative PE condition [39]. Such end-to-end insights are a valuable feature of direct data-driven control methods, which are often harder to achieve in indirect approaches. While the original paper [51] required an additional non-convex constraint $\lVert\mathbf{\sigma}(t)\rVert_{\infty}\leq\bar{\mathbf{\varepsilon}}(1+\lVert\mathbf{\alpha}(t)\rVert_{1})$ to bound the slack variable, it was later shown in [58] that this constraint can be dropped when $\lambda_{\mathbf{\sigma}}\sim\frac{1}{\bar{\mathbf{\varepsilon}}}$, see also [70, Theorem 4.1].
@@
As in the noise-free case, the occurrence of terminal equality constraints poses practical limitations, e.g., a small region of attraction, and alternative means of guaranteeing closed-loop stability are available. Roughly speaking, any model-based MPC scheme which admits inherent robustness properties against disturbances can be transformed into a data-driven MPC scheme with closed-loop guarantees in the presence of noise [71]. Beyond this general principle, the literature contains various results on closed-loop guarantees in robust data-driven MPC. For example, in the noisy data case, terminal ingredients can be designed based on robust data-driven output-feedback design [55]. By combining continuity of data-driven MPC with respect to noise [72] and inherent robustness properties [71, 73], one can show that the resulting data-driven MPC scheme admits closed-loop guarantees also in the presence of noise, compare [70, Section 4.5.2]. Likewise, data-driven MPC without terminal ingredients practically exponentially stabilizes the closed-loop system, assuming that the prediction horizon $L$ is sufficiently long [58]. Notably, in both cases (terminal ingredients and $L$ sufficiently long), stability and robustness can be proven even for a $1$-step MPC scheme as in Algorithm 4.1. Alternatively, one can optimize over an artifical equilibrium online which provides powerful practical features (no knowledge required if $(\mathbf{u}^{\mathrm{s}},\mathbf{y}^{\mathrm{s}})$ is a feasible equilibrium, improved robustness, increased region of attraction, handling online setpoint changes) and also admits closed-loop guarantees in the presence of noise [72].